By De Morgan's Law, Ec∩Fc=(E∪F)c. Therefore:
P(Ec∩Fc∣G)=1−P(E∪F∣G)
By the addition theorem of probability conditioned on G:
P(E∪F∣G)=P(E∣G)+P(F∣G)−P(E∩F∣G)
-
Since E,G and F,G are pairwise independent:
P(E∣G)=P(E),P(F∣G)=P(F)
-
The conditional joint probability is:
P(E∩F∣G)=P(G)P(E∩F∩G)=P(G)0=0
Substituting these back:
P(E∪F∣G)=P(E)+P(F)−0=P(E)+P(F)
Therefore:
P(Ec∩Fc∣G)=1−[P(E)+P(F)]=[1−P(E)]−P(F)=P(Ec)−P(F)
Correct Answer: C